OpenStudy (anonymous):

Solve the equation in quadratic form. 4x^4=13x^2-9

OpenStudy (anonymous):

I did the first step and got 4x^4-13x^2+9=0 Then I factored by grouping and got [4x^4-4x][-9x+9] and im stuck on what to do next.

OpenStudy (anonymous):

Set both factors equal to zero and solve for x. Make sure to check your answers.

OpenStudy (anonymous):

Even with variables left?

OpenStudy (anonymous):

Yeah -9x+9=0 4x^4-4x=0

OpenStudy (anonymous):

You can factor out fours and nines to make it easier

OpenStudy (anonymous):

so i just take out 4x and 9? because thats what I did before and I got 4x(x and I dont know what else is supposed to be in the ( )

OpenStudy (anonymous):

Once you factor out, you can divide both sides of thes equation by your factors, which eliminates them from consequence in the equation\[x-1=0\]\[ x ^{3}-1=0\]

OpenStudy (anonymous):

so did you divide by four somewhere in there

OpenStudy (anonymous):

the 4x(x-1) maybe?

OpenStudy (anonymous):

No, I only divided the bottom equation by 4x

OpenStudy (anonymous):

Oh well in the end did u get